OpenStudy (anonymous):

I don't understand a question on my homework.. Over the past 7 years the median rainfall in West Berry has been 74 inches. The greatest rainfall was 102 inches. The least was 52 inches. List possible rainfall amounts for the 7 years. I can't focus and I really need help.

OpenStudy (texaschic101):

Over the past 7 years, the median (average) has been 74 inches you are given 2 amounts of rainfall --- 102 and 52 (102 + 52 + 5x)/7 = 74 (5x + 154)/7 = 74 5x + 154 = 7 * 74 5x + 154 = 518 5x = 518 - 154 5x = 364 x = 364/5 x = 72.8 check.. ((102 + 52 + 5(72.8))/7 = 74 (102 + 52 + 364) / 7 = 74 518/7 = 74 74 = 74 (correct) so one year, rainfall was 52 inches....another year, rainfall was 102 inches....and for the other 5 years, it was 72.8 inches per year...thus, averaging 74 inches. you might better get a second opinion on this

OpenStudy (texaschic101):

@tkhunny ......The more I look at this problem, the more I doubt my answer...can you take a peek at this.....am I totally wrong on this one ??

OpenStudy (texaschic101):

or does this problem require a box and whisker plot

OpenStudy (tkhunny):

You get 7 years. 52 and 102 are chose for you. There are only 5 more to select. Beings as there are five (odd number) more to select, the Median, 74 must also be selected, leaving only four more. Note: I'm assuming there are no repeats. Thus: 52 __ __ 74 __ __ 102 There are quite a few ways to fill in the four blanks.

OpenStudy (texaschic101):

ohhhh....I see....I was way off. Thanks so much for explaining :)